harrys motorboat can make an average of 8 miles per hour . one day he sets out for a trip , only to have the motor break down . harry rows back at two miles an hour. When he reaches his dock , he finds that he has been gone 5 hours . how far has he rowed?
Answer by ankor@dixie-net.com(22740)

Harry's motorboat can make an average of 8 miles per hour. One day he sets out for a trip, only to have the motor break down. Harry rows back at two miles an hour. When he reaches his dock, he finds that he has been gone 5 hours. how far has he rowed?
:
Let d = distance he rowed, (also the distance he went by motor)
:
Write a time equation: Time = Dist/speed
:
Motor time + Rowing time = 5 hrs
d/8 + d/2 = 5
:
multiply equation by 8:
d + 4d = 40
5d = 40
d = 40/5
d = 8 miles
:
Check solution 8/8 + 8/2 = 5 hrs